7. A firm is considering raising its price by 9 percent and has hired an econometrician to estimate the elasticity of demand for its product. The econometrician estimates the parameters of a log-liner demand function and reports that the parameter estimate for the elasticity of demand is -1.5 and the standard error of the estimate is 0.3.
a. If the firm raises its price by 9 percent, what is the expected change in quantity demanded?
b. Approximate the upper and lower bounds on the 95 percent confidence interval for the change in quantity demanded.
Answer #1

7.

A

Elasticity of demand = % change in quantity demanded / % change in Price

% change in quantity demanded = -1.5*9% = -13.5%

So,

Quantity demanded will decrease by 13.5%.

B.

At 95% confidence interval,

Lower bound = -1.5 - 2*(.3) = -2.1

Upper bound = -1.5 + 2*(.3) = -.9

At lower bound, decrease in quantity demanded = 2.1*9% = 18.9%

At upper bound, decrease in quantity demanded = .9*9% = 8.1%